# Sequential Kernelized Stein Discrepancy

### v1

@@ Abstract
We present a sequential version of the kernelized Stein discrepancy, which allows for conducting goodness-of-fit tests for unnormalized densities that are continuously monitored and adaptively stopped. That is, the sample size need not be fixed prior to data collection; the practitioner can choose whether to stop the test or continue to gather evidence at any time while controlling the false discovery rate. In stark contrast to related literature, we do not impose uniform boundedness on the Stein kernel. Instead, we exploit the potential boundedness of the Stein kernel at arbitrary point evaluations to define test martingales, that give way to the subsequent novel sequential tests. We prove the validity of the test, as well as an asymptotic lower bound for the logarithmic growth of the wealth process under the alternative. We further illustrate the empirical performance of the test with a variety of distributions, including restricted Boltzmann machines.
@@ 1 Introduction
Many statistical procedures heavily rely on the assumptions made about the distribution of the empirical observations, and prove invalid if such conditions are violated. For instance, a high-energy physicist may develop a generative model seeking to produce synthetic observations of particle collisions, motivated by the high cost of experimenting on a real particle collider. If the model is not accurate, then any conclusion that derives from such synthetic data will lack any scientific interest. The question of whether the data follows a particular distribution may be posed in terms of goodness-of-fit testing. Formally, the simple goodness-of-fit testing problem considers the null hypothesis $H_{0}:Q=P$ against the alternative $H_{1}:Q\neq P$, for a given $P$ and access to data $X_{1},X_{2},\ldots\sim Q$. It may also be the case that we have a set of given distributions as candidates and we wish to test if any of them accurately models the empirical data. In such a case, the goodness-of-fit problem is posed in terms of a composite null hypothesis $H_{0}:Q\in\mathcal{P}$, against the alternative hypothesis $H_{1}:Q\not\in\mathcal{P}$.
@@
In this work, we focus on a particularly challenging scenario that arises from not having full access to $P$ (or $\mathcal{P}$). We handle classes of distributions whose densities are only known up to normalizing constants. This is a common scenario when working with general energy-based models, such as Ising models and restricted Boltzmann machines, and in Bayesian statistics, where normalizing factors are generally neither known in closed form nor computable. The existing works in the literature regarding goodness-of-fit for unnormalized densities have focused on the batch or fixed sample size setting, where the number of samples $n$ is decided before conducting the analysis, which is later on carried on using observations $X_{1},\ldots,X_{n}$. Nonetheless, this approach counts with several major drawbacks. Because the number of observations in batch tests is determined prior to data collection, there is a risk that too many observations may be allocated to simpler problem instances, thus wasting resources, or that insufficient observations are assigned to more complex instances, which can yield insufficient evidence against the null hypothesis. Furthermore, when test outcomes appear encouraging but not definitive (for instance, if a p-value is marginally higher than a specified significance threshold), one may be tempted to augment the dataset and undertake the investigation again. Traditional batch testing methods, however, do not allow for this approach.
@@
In contrast, we seek to develop tests that are anytime valid. That is, we can repeatedly decide whether to collect more data based on the current state of the procedure without compromising later assessments, halting the process at any time and for any reason. Formally, a sequential test which is stopped at any given arbitrary moment can be represented by a random stopping time $\tau$ taking values in $\{1,2,\ldots\}\cup\{\infty\}$. The stopping time $\tau$ denotes the random time at which the null hypothesis is rejected. A sequential test is called level-$\alpha$ if it satisfies the condition $\mathbb{P}(\tau<\infty)\leq\alpha$ under the null for any stopping time $\tau$.
@@
In this work, we develop level-$\alpha$ sequential goodness-of-fit tests that accommodate for unnormalized densities, building on the concept of kernel Stein discrepancies. The type I error is controlled even if the tests are continuously monitored and adaptively stopped, hence automatically adapting the sample size to the unknown alternative. In particular, we believe these anytime-valid tests will be of remarkable interest in the following scenarios:
@@
  • Measuring the quality of a proposed distribution: We may have a specific candidate distribution (or set of candidate distributions) to model the data that does not count with a computable normalizing constant (e.g., some complex network that models the interaction between genes). Does this distribution accurately model the true data generation process?
@@
  • Measuring sample quality: We may have obtained an unnormalized density, such as some posterior distribution in Bayesian statistics, that we wish to draw from using Markov Chain Monte Carlo (MCMC) methods. Is the chosen MCMC algorithm yielding samples accurately from such a density?
@@
The work is organized as follows. Section 2 presents the related work. Section 3 introduces preliminary work, with special emphasis on the kernel Stein discrepancy and testing by betting. These constitute the theoretical foundations of the novel goodness-of-fit tests, presented in Section 4. Section 5 subsequently provides a general derivation for the lower bounds required in the test, alongside different examples. The empirical validity of the tests is presented in Section 6, followed by concluding remarks in Section 7.
@@ 2 Related work
Concurrently, kernel methods have received widespread attention due to their notable empirical performance and theoretical guarantees. Reproducing kernel Hilbert spaces provide the theoretical foundation for these methods (Schölkopf and Smola,, 2002; Berlinet and Thomas-Agnan,, 2011). We highlight three predominant applications of kernel methods. First, the maximum mean discrepancy (MMD) has given way to kernel-based two sample tests (Gretton et al.,, 2006, 2012). Second, independence tests have been developed based on the Hilbert Schmidt independence criterion (HSIC) (Gretton et al.,, 2005, 2007). Third, kernel Stein discrepancies (KSD) have led to novel goodness-of-fit tests (Liu et al.,, 2016; Chwialkowski et al.,, 2016; Gorham and Mackey,, 2017).
At the intersection of testing by betting and kernels, Shekhar and Ramdas, (2023) developed a sequential MMD for conducting two sample tests that can be arbitrarily stopped. Similarly, Podkopaev et al., (2023) extended the HSIC to the sequential setting. Both works rely on the uniform boundedness of the chosen kernel, which is key for the construction of nonnegative martingales. While some kernels commonly used for the MMD and HSIC (such as the RBF or Laplace kernels) are uniformly bounded, this is very rarely the case for the Stein kernel exploited by the KSD. The methodology proposed in this contribution does not assume uniform boundedness of the Stein kernel, which poses a number of theoretical challenges that will be addressed in the subsequent sections.
@@ 3 Background
3.1 The kernel Stein discrepancy
Throughout this contribution we will draw upon the kernel Stein discrepancy (KSD), which is a distributional divergence that builds on the concept of reproducing kernel Hilbert space (RKHS) and the general Stein’s method (Stein,, 1972; Gorham and Mackey,, 2015). The KSD may be understood as a kernelized version of score-matching divergence (Hyvärinen and Dayan,, 2005).
@@ 3.1 The kernel Stein discrepancy
Kernel Stein discrepancies (KSD): Assume now that $P$ has density $p$ on $\mathcal{X}\subset\mathbb{R}^{d}$, and let $\mathcal{H}$ be an RKHS with reproducing kernel $k$ such that $\nabla k(\cdot,x)$ exists for all $x\in\mathcal{X}$. Based on the Stein operator
$(T_{p}h)(x)=\sum_{i=1}^{d}\left(\frac{\partial\log{p}(x)}{\partial x_{i}}h_{i}(x)+\frac{\partial h_{i}(x)}{\partial x_{i}}\right),\quad h\in\mathcal{H}^{d},$
@@
$\text{KSD}_{\mathcal{H}}(Q,P)=\sup_{f\in\mathcal{F}_{\text{MMD}}}\mathbb{E}_{X\sim Q}[(T_{p}f)(X)],$
where $\mathcal{F}_{\text{KSD}}=\{h\in\mathcal{H}^{d}:\|h\|_{\mathcal{H}^{d}}\leq 1\}$. Defining $s_{p}(x):=\nabla_{x}\log p(x)$ and $\xi_{p}(\cdot,x):=\left[s_{p}(x)k(\cdot,x)+\nabla k(\cdot,x)\right]\in\mathcal{H}^{d}$, it follows that
@@
$=\langle s_{p}(x),s_{p}(\tilde{x})\rangle_{\mathbb{R}^{d}}k(x,\tilde{x})+\langle s_{p}(\tilde{x}),\nabla_{x}k(x,\tilde{x})\rangle_{\mathbb{R}^{d}}$
@@
is a reproducing kernel based on the Moore-Aronszajn theorem. Again, if $\mathbb{E}_{X\sim Q}\sqrt{h_{p}(X,X)}<\infty$, then $\text{KSD}_{\mathcal{H}}(Q,P)=\|\mathbb{E}_{X\sim Q}\left[\xi_{p}(\cdot,X)\right]\|_{\mathcal{H}^{d}}$. If $k$ is universal and under certain regularity conditions, then $\text{KSD}_{\mathcal{H}}(Q,P)=0$ if, and only if, $Q=P$ (Chwialkowski et al.,, 2016). In the batch setting and simple null hyposthesis, a test statistic is generally taken as a V-statistic or U-statistic, and parametric or wild bootstrap is used to calibrate the test. We refer the reader to Key et al., (2021) for the more challenging composite null hypothesis case.
@@ 4 Sequential goodness-of-fit by betting
$\mathbb{E}_{H_{0}}\left[\frac{1}{t-1}\sum_{i=1}^{t-1}h_{p}(X_{i},X_{t})\bigg{|}\mathcal{F}_{t-1}\right]$ $=\mathbb{E}_{H_{0}}\left[\frac{1}{t-1}\sum_{i=1}^{t-1}\langle\xi_{p}(\cdot,X_{i}),\xi_{p}(\cdot,X_{t})\rangle_{\mathcal{H}^{d}}\bigg{|}\mathcal{F}_{t-1}\right]$ (4)
@@
$h_{p}(x,\tilde{x})=\langle\xi_{p}(\cdot,x),\xi_{p}(\cdot,\tilde{x})\rangle_{\mathcal{H}^{d}}=\|\xi_{p}(\cdot,x)\|_{\mathcal{H}^{d}}\|\xi_{p}(\cdot,\tilde{x})\|_{\mathcal{H}^{d}}\cos\beta(x,\tilde{x}),$
@@ 4.1 Simple null hypothesis
is lower bounded by -1. Hence, as long as the betting strategy $\lambda_{t}$ is nonnegative and bounded above by 1, we are able to define a wealth process that forms a test martingale. The following theorem formally establishes such a result; the proof is deferred to Appendix B.
@@
Assume that $\mathbb{E}[h_{p}(X,X^{\prime})]=0$, and let $\lambda_{t}\in[0,1]$ be predictable. The wealth process
@@
Algorithm 1 summarizes the procedure introduced in this section, with a complexity of $O(T^{2})$ for $T$ rounds. Note that $\mathbb{E}[h_{p}(X,X^{\prime})]=0$ under the null if $\mathbb{E}[h_{p}(X,X)]<\infty$, and thus in that case Theorem 1 establishes the validity of Algorithm 1 for any betting strategy that is predictable. However, the power of the test will heavily depend on the chosen betting strategy. For instance, if we take $\lambda_{t}=0$ for all $t$ (this is, we never bet any money), then the wealth will remain constant as $\mathcal{K}_{t}=1$, and so the test is powerless. There exists a variety of betting strategies that have been studied in the literature. In this contribution, we focus on aGRAPA (‘approximate GRAPA’) and LBOW (‘Lower-Bound On the Wealth’). We refer the reader to Waudby-Smith and Ramdas, (2024) to a detailed presentation of different betting strategies.
@@
$\lambda_{t+1}^{\text{aGRAPA}}=\min\left(1,\max\left(0,\frac{\frac{1}{t-1}\sum_{i=1}^{t-1}g_{i}(X_{i})}{\frac{1}{t-1}\sum_{i=1}^{t-1}g_{i}^{2}(X_{i})}\right)\right).$ (7)
@@
$\lambda_{t+1}^{\text{LBOW}}=\max\left(0,\frac{\frac{1}{t-1}\sum_{i=1}^{t-1}g_{i}(X_{i})}{\frac{1}{t-1}\sum_{i=1}^{t-1}g_{i}(X_{i})+\frac{1}{t-1}\sum_{i=1}^{t-1}g_{i}^{2}(X_{i})}\right).$ (8)
@@
Assume $\mathbb{E}[h_{p}(X,X)]<\infty$ and $\mathbb{E}[M_{p}(X)]<\infty$. Denote $g^{*}(x):=\mathbb{E}[h_{p}(X,x)]/\mathbb{E}[M_{p}(X)]$. Under $H_{1}$, if $\mathbb{E}[g^{*}(X)]>0$, the LBOW betting strategy yields
@@
The proof of Theorem 2 is deferred to Appendix B. We point out that the unboundedness of the Stein kernel does not allow for easily extending the arguments presented in Podkopaev et al., (2023), which rely on a different betting strategy whose guarantees stem from uniformly bounded payoff functions. Furthermore, we highlight the mildness of the assumptions in Theorem 2. Assumption $\mathbb{E}[M_{p}(X)]<\infty$ only requires the first moment of the upper bounds to exist. This condition usually reduces to the existence of a specific moment of the original distribution, which is often easily verifiable. Assumption $\mathbb{E}[h_{p}(X,X)]<\infty$ is an ubiquitous assumption in the KSD theory; it is equivalent to the existence of the second moment of $\|\xi_{p}(\cdot,X)\|_{\mathcal{H}^{d}}$, which is precisely the theoretical object that the KSD builds on. Finally, note that $P\neq Q$ does not necessarily imply $\mathbb{E}[g^{*}(X)]>0$ unless some regularity conditions hold, such as $C_{0}$-universality of $k$ and $\mathbb{E}[\nabla\log(p(X)/q(X))]<\infty$. These conditions have been extensively studied in the batch setting (Chwialkowski et al.,, 2016; Liu et al.,, 2016; Gorham and Mackey,, 2017), and are inherent to the KSD approach.

### v2

@@ Abstract
We present a sequential version of the kernelized Stein discrepancy goodness-of-fit test, which allows for conducting goodness-of-fit tests for unnormalized densities that are continuously monitored and adaptively stopped. That is, the sample size need not be fixed prior to data collection; the practitioner can choose whether to stop the test or continue to gather evidence at any time while controlling the false discovery rate. In stark contrast to related literature, we do not impose uniform boundedness on the Stein kernel. Instead, we exploit the potential boundedness of the Stein kernel at arbitrary point evaluations to define test martingales, that give way to the subsequent novel sequential tests. We prove the validity of the test, as well as an asymptotic lower bound for the logarithmic growth of the wealth process under the alternative. We further illustrate the empirical performance of the test with a variety of distributions, including restricted Boltzmann machines.
@@ 1 INTRODUCTION
Many statistical procedures heavily rely on the assumptions made about the distribution of the empirical observations, and prove invalid if such conditions are violated. For instance, a high-energy physicist may develop a generative model seeking to produce synthetic observations of particle collisions, motivated by the high cost of experimenting on a real particle collider (Agostinelli et al.,, 2003; Chekalina et al.,, 2019). If the model is not accurate, then any conclusion that derives from such synthetic data will lack any scientific interest (Huang et al.,, 2023; Masserano et al.,, 2023). The question of whether the data follows a particular distribution may be posed in terms of goodness-of-fit testing (Lehmann et al.,, 1986; González-Manteiga and Crujeiras,, 2011; D’Agostino,, 2017). Formally, the simple goodness-of-fit testing problem considers the null hypothesis $H_{0}:Q=P$ against the alternative $H_{1}:Q\neq P$, for a given $P$ and access to data $X_{1},X_{2},\ldots\sim Q$. It may also be the case that we have a set of given distributions as candidates and we wish to test if any of them accurately models the empirical data (Durbin,, 1975; Key et al.,, 2021). In such a case, the goodness-of-fit problem is posed in terms of a composite null hypothesis $H_{0}:Q\in\mathcal{P}$, against the alternative hypothesis $H_{1}:Q\not\in\mathcal{P}$.
@@
In this work, we focus on a particularly challenging scenario that arises from not having full access to $P$ (or $\mathcal{P}$). We handle classes of distributions whose densities are only known up to normalizing constants. This is a common scenario when working with general energy-based models (LeCun et al.,, 2006; Du and Mordatch,, 2019), such as Ising models (Geman and Geman,, 1984; Clifford,, 1990) and restricted Boltzmann machines (Ackley et al.,, 1985; Hinton,, 2010), and in Bayesian statistics, where normalizing factors are generally neither known in closed form nor computable (Bolstad and Curran,, 2016). Most of the existing works in the literature regarding goodness-of-fit for unnormalized densities have focused on the batch or fixed sample size setting, where the number of samples $n$ is decided before conducting the analysis, which is later on carried on using observations $X_{1},\ldots,X_{n}$ (Liu et al.,, 2016; Chwialkowski et al.,, 2016; Gorham and Mackey,, 2017). Nonetheless, this approach comes with several major drawbacks. Because the number of observations in batch tests is determined prior to data collection, there is a risk that too many observations may be allocated to simpler problem instances, thus wasting resources, or that insufficient observations are assigned to more complex instances, which can yield insufficient evidence against the null hypothesis. Furthermore, when test outcomes appear encouraging but not definitive (for instance, if a p-value is marginally higher than a specified significance threshold), one may be tempted to augment the dataset and undertake the investigation again. Traditional batch testing methods, however, do not allow for this approach.
@@
In contrast, we seek to develop tests that are anytime valid (Ramdas et al.,, 2023; Grünwald et al.,, 2024). That is, we can repeatedly decide whether to collect more data based on the current state of the procedure without compromising later assessments, halting the process at any time and for any reason. Formally, a sequential test which is stopped at any given arbitrary moment can be represented by a random stopping time $\tau$ taking values in $\{1,2,\ldots\}\cup\{\infty\}$. The stopping time $\tau$ denotes the random time at which the null hypothesis is rejected. A sequential test is called level-$\alpha$ if it satisfies the condition $\mathbb{P}(\tau<\infty)\leq\alpha$ under the null for any stopping time $\tau$.
@@
In this work, we develop level-$\alpha$ sequential goodness-of-fit tests that accommodate for unnormalized densities, building on the concept of kernelized Stein discrepancies. The type I error is controlled even if the tests are continuously monitored and adaptively stopped, hence automatically adapting the sample size to the unknown alternative. In particular, we believe these anytime-valid tests will be of remarkable interest in the following scenarios:
@@
  • Measuring the quality of a proposed distribution: We may have a specific candidate distribution (or set of candidate distributions) to model the data whose normalizing constant is unknown (e.g., some complex network that models the interaction between genes). Does this distribution accurately model the true data generation process?
@@
  • Measuring sample quality: We may have obtained an unnormalized density, such as some posterior distribution in Bayesian statistics, that we wish to draw from using Markov Chain Monte Carlo (MCMC) methods. Is the chosen MCMC algorithm yielding samples accurately from such a density? [^1] There exist complexities when sequentially analyzing the fitness of MCMC samples, mainly related to autocorrelation and burn-in stages. In order to minimize the effect of autocorrelation, one could skip samples and only input one every X samples for the sequential test, where X is large enough to considerably lower the correlation between samples. In order to avoid early rejection due to burn-in stage, sequential tests can be initialized after such a burn-in period.
@@
The work is organized as follows. Section 2 presents the related work. Section 3 introduces preliminary work, with special emphasis on the kernelized Stein discrepancy and testing by betting. These constitute the theoretical foundations of the novel goodness-of-fit tests, presented in Section 4. Section 5 subsequently provides a general derivation for the lower bounds required in the test, alongside different examples. The empirical validity of the tests is presented in Section 6, followed by concluding remarks in Section 7.
@@ 2 RELATED WORK
Concurrently, kernel methods have received widespread attention due to their notable empirical performance and theoretical guarantees. Reproducing kernel Hilbert spaces provide the theoretical foundation for these methods (Schölkopf and Smola,, 2002; Berlinet and Thomas-Agnan,, 2011). We highlight three predominant applications of kernel methods. First, the maximum mean discrepancy (MMD) has given way to kernel-based two sample tests (Gretton et al.,, 2006, 2012). Second, independence tests have been developed based on the Hilbert Schmidt independence criterion (HSIC) (Gretton et al.,, 2005, 2007). Third, kernelized Stein discrepancies (KSD) have led to novel goodness-of-fit tests (Liu et al.,, 2016; Chwialkowski et al.,, 2016; Gorham and Mackey,, 2017). While the analysis of this contribution is tailored to so-called Langevin KSD, the ideas presented herein could potentially be extended to abstract domains including categorical data (Yang et al.,, 2018), censored data (Fernandez et al.,, 2020), directional data (Xu and Matsuda,, 2020), functional data (Wynne et al.,, 2025), sequential data (Baum et al.,, 2023), and point processes (Yang et al.,, 2019).
At the intersection of testing by betting and kernels, Shekhar and Ramdas, (2023) developed a sequential MMD for conducting two sample tests that can be arbitrarily stopped. Similarly, Podkopaev et al., (2023) extended the HSIC to the sequential setting. Recently, Zhou and Liu, (2024) proposed a sequential KSD, in a similar spirit to our contribution. [^2] We defer a comprehensive comparison of Zhou and Liu, (2024) and our contribution to Appendix B. These three works rely on the uniform boundedness of the chosen kernel, which is key for the construction of nonnegative martingales. While some kernels commonly used for the MMD and HSIC (such as the RBF or Laplace kernels) are uniformly bounded, this is very rarely the case for the Stein kernel exploited by the KSD. The methodology proposed in this contribution does not assume uniform boundedness of the Stein kernel, which poses a number of theoretical challenges that will be addressed in the subsequent sections.
@@ 3 BACKGROUND
3.1 The Kernelized Stein Discrepancy
Throughout this contribution we will draw upon the kernelized Stein discrepancy (KSD), which is a distributional divergence that builds on the concept of reproducing kernel Hilbert space (RKHS) and the general Stein’s method (Stein,, 1972; Gorham and Mackey,, 2015). The KSD may be understood as a kernelized version of score-matching divergence (Hyvärinen and Dayan,, 2005).
@@ 3.1 The Kernelized Stein Discrepancy
Kernelized Stein discrepancies (KSD): Assume now that $P$ has density $p$ on $\mathcal{X}\subset\mathbb{R}^{d}$, and let $\mathcal{H}$ be an RKHS with reproducing kernel $k$ such that $\nabla k(\cdot,x)$ exists for all $x\in\mathcal{X}$. Based on the Stein operator
$(T_{p}h)(x)=\sum_{i=1}^{d}\left(\frac{\partial\log{p}(x)}{\partial x_{i}}h_{i}(x)+\frac{\partial h_{i}(x)}{\partial x_{i}}\right),\;h\in\mathcal{H}^{d},$
@@
$\text{KSD}_{\mathcal{H}}(Q,P)=\sup_{f\in\mathcal{F}_{\text{KSD}}}\mathbb{E}_{X\sim Q}[(T_{p}f)(X)],$
where we consider the set $\mathcal{F}_{\text{KSD}}=\{h\in\mathcal{H}^{d}:\|h\|_{\mathcal{H}^{d}}\leq 1\}$. Defining $s_{p}(x):=\nabla_{x}\log p(x)$ and $\xi_{p}(\cdot,x):=\left[s_{p}(x)k(\cdot,x)+\nabla k(\cdot,x)\right]\in\mathcal{H}^{d}$, it follows that
@@
$=\langle s_{p}(x),s_{p}(\tilde{x})\rangle_{\mathbb{R}^{d}}k(x,\tilde{x})$
$\quad+\langle s_{p}(\tilde{x}),\nabla_{x}k(x,\tilde{x})\rangle_{\mathbb{R}^{d}}$
@@
is a reproducing kernel based on the Moore-Aronszajn theorem. Again, if $\mathbb{E}_{X\sim Q}\sqrt{h_{p}(X,X)}<\infty$, then $\text{KSD}_{\mathcal{H}}(Q,P)=\|\mathbb{E}_{X\sim Q}\left[\xi_{p}(\cdot,X)\right]\|_{\mathcal{H}^{d}}$. If $k$ is universal and under certain regularity conditions, then $\text{KSD}_{\mathcal{H}}(Q,P)=0$ if, and only if, $Q=P$ (Chwialkowski et al.,, 2016). In the batch setting and simple null hyposthesis, a test statistic is generally taken as a V-statistic or U-statistic, and parametric or wild bootstrap is used to calibrate the test. We refer the reader to Key et al., (2021) for the more challenging composite null hypothesis case. [^3] In the context of approximating an integral, the KSD also allows for evaluating the realized empirical measure (rather than the marginal distribution). We refer the reader to Kanagawa et al., (2022) for a recent account of this research direction.
@@ 4 SEQUENTIAL GOODNESS-OF-FIT BY BETTING
$\mathbb{E}_{H_{0}}\left[\frac{1}{t-1}\sum_{i=1}^{t-1}h_{p}(X_{i},X_{t})\bigg{|}\mathcal{F}_{t-1}\right]$ (4)
$=$ $\mathbb{E}_{H_{0}}\left[\frac{1}{t-1}\sum_{i=1}^{t-1}\langle\xi_{p}(\cdot,X_{i}),\xi_{p}(\cdot,X_{t})\rangle_{\mathcal{H}^{d}}\bigg{|}\mathcal{F}_{t-1}\right]$
@@
$h_{p}(x,\tilde{x})$ $=\langle\xi_{p}(\cdot,x),\xi_{p}(\cdot,\tilde{x})\rangle_{\mathcal{H}^{d}}$
$=\|\xi_{p}(\cdot,x)\|_{\mathcal{H}^{d}}\|\xi_{p}(\cdot,\tilde{x})\|_{\mathcal{H}^{d}}\cos\beta(x,\tilde{x}),$
@@ 4.1 Simple Null Hypothesis
is lower bounded by -1. Hence, as long as the betting strategy $\lambda_{t}$ is nonnegative and bounded above by 1, we are able to define a wealth process that forms a test martingale. The following theorem formally establishes such a result; the proof is deferred to Appendix C.
@@
Assume that $\mathbb{E}_{H_{0}}[h_{p}(X,X^{\prime})]=0$, and let $\lambda_{t}\in[0,1]$ be predictable. The wealth process
@@
Algorithm 1 summarizes the procedure introduced in this section. Given that the complexity of computing $g_{t}$ is $O(t)$ for each $t$, the complexity of Algorithm 1 for $T$ rounds is $O(T^{2})$. Note that $\mathbb{E}[h_{p}(X,X^{\prime})]=0$ under the null if $\mathbb{E}[h_{p}(X,X)]<\infty$, and thus in that case Theorem 1 establishes the validity of Algorithm 1 for any betting strategy that is predictable. However, the power of the test will heavily depend on the chosen betting strategy. For instance, if we take $\lambda_{t}=0$ for all $t$ (this is, we never bet any money), then the wealth will remain constant as $\mathcal{K}_{t}=1$, and so the test is powerless. There exists a variety of betting strategies that have been studied in the literature. In this contribution, we focus on aGRAPA (‘approximate GRAPA’) and LBOW (‘Lower-Bound On the Wealth’). We refer the reader to Waudby-Smith and Ramdas, (2024) to a detailed presentation of different betting strategies.
@@
$\lambda_{t+1}^{\text{aGRAPA}}=1\wedge\left(0\vee\frac{\frac{1}{t-1}\sum_{i=1}^{t-1}g_{i}(X_{i})}{\frac{1}{t-1}\sum_{i=1}^{t-1}g_{i}^{2}(X_{i})}\right).$ (7)
@@
$\lambda_{t+1}^{\text{LBOW}}=0\vee\frac{\frac{1}{t-1}\sum_{i=1}^{t-1}g_{i}(X_{i})}{\frac{1}{t-1}\sum_{i=1}^{t-1}g_{i}(X_{i})+\frac{1}{t-1}\sum_{i=1}^{t-1}g_{i}^{2}(X_{i})}.$ (8)
@@
Let $(X_{i})_{i\geq 1}$ be independent and identically distributed copies of $X$. Assume $\mathbb{E}[h_{p}(X,X)]<\infty$ and $\mathbb{E}[M_{p}(X)]<\infty$. Denote $g^{*}(x):=\mathbb{E}[h_{p}(X,x)]/\mathbb{E}[M_{p}(X)]$. Under $H_{1}$, if $\mathbb{E}[g^{*}(X)]>0$, the LBOW betting strategy yields
@@
The proof of Theorem 2 is deferred to Appendix C. We point out that the unboundedness of the Stein kernel does not allow for easily extending the arguments presented in Podkopaev et al., (2023), which rely on a different betting strategy whose guarantees stem from uniformly bounded payoff functions. Furthermore, we highlight the mildness of the assumptions in Theorem 2. Assumption $\mathbb{E}[M_{p}(X)]<\infty$ only requires the first moment of the upper bounds to exist. This condition usually reduces to the existence of a specific moment of the original distribution, which is often easily verifiable. Assumption $\mathbb{E}[h_{p}(X,X)]<\infty$ is an ubiquitous assumption in the KSD theory; it is equivalent to the existence of the second moment of $\|\xi_{p}(\cdot,X)\|_{\mathcal{H}^{d}}$, which is precisely the theoretical object that the KSD builds on. Finally, note that $P\neq Q$ does not necessarily imply $\mathbb{E}[g^{*}(X)]>0$. However, there exist sufficient conditions for this implication to hold, such as $C_{0}$-universality of $k$ and a moment condition on $\nabla\log(p(X)/q(X)$ (Chwialkowski et al.,, 2016, Theorem 2.2.) or a root exponential growth of $\|s_{p}\|$ for a rich family of base kernels (Barp et al.,, 2024, Application 2). These conditions have already been extensively studied in the batch setting (Chwialkowski et al.,, 2016; Liu et al.,, 2016; Gorham and Mackey,, 2017; Barp et al.,, 2024), and equally apply to our setting.